# The Pólya-Tchebotarev problem with semiclassical external fields

### v1

@@ 1 Introduction
A classical problem in Geometric Function Theory, known as the Pólya-Chebotarev Problem, asks for finding a compact and connected set of minimal capacity that contains a given finite collection of points $\mathcal{C}\subset\mathbb{C}$ [23, 22]. In terms of the closely related notion of logarithmic energy, the Pólya-Chebotarev Problem may be phrased as the question of finding a compact and connected set $K_{*}\subset\mathbb{C}$ containing $\mathcal{C}$, and for which the max-min (or sup-inf) problem
@@
The original form of the problem was presented by Pólya in 1929 [23], and solved by Lavrentieff [16, 15], Grötzch [13] and Goluzin [11] in the following years. Much later, motivated by questions on Padé approximants and rational approximation [33, 29, 21], Stahl established a more general version of Pólya-Chebotarev problem [31, 30], in that it weakens the finiteness condition on $\mathcal{C}$, allowing it to be an arbitrary compact set with zero capacity.
Stahl’s ideas were rather influential in particular in the theory of orthogonal polynomials. In his developments, he also connected sets of minimal capacity with non-hermitian orthogonal polynomials, showing that in the large degree limit the zeros of such polynomials are described by sets of minimal capacity [32], see also [12]. Starting in the early 2000s, such ideas and results also found their way with applications to other areas of mathematics such as numerical analysis [10, 8, 9], random matrix theory [1, 2, 6, 3], rational solutions to integrable systems [7, 5], among others.
In many of these applications we just mentioned, the corresponding extremal problems that arise often are analogues of the original Pólya-Chebotarev Problem, but now under the setting of weighted capacity, which is the main focus of our work here. We solve a weighted extension of the original the Pólya-Chebotarev Problem (1.1), obtained including an external field in the energy functional: for a given finite set $\mathcal{C}\subset\mathbb{C}$ and a suitable function $\varphi$ defined over $\mathbb{C}$, find a compact and connected set $K_{*}\subset\mathbb{C}$ containing $\mathcal{C}$, and for which the max-min problem
@@
When $\Phi$ is a constant/linear/quadratic real-valued polynomial and $\mathcal{T}=\mathcal{T}_{\text{Jac}}/\mathcal{T}_{\text{Lag}}/\mathcal{T}_{\text{Herm}}$, the polynomial $p_{n}$ reduces to (particular cases of rescaled) Jacobi/Laguerre/Hermite polynomials, respectively. For general complex-valued polynomial $\Phi$, the polynomial $p_{n}$, if it exists, is independent of the contour of orthogonality $\Gamma\in\mathcal{T}$ chosen, since the orthogonality condition (1.3) remains unchanged if we deform $\Gamma$ to a new contour $\widetilde{\Gamma}$ within the same class.
In each of the aforementioned cases, the flexibility in the choice of contours $\Gamma\in\mathcal{T}$ has impact in the behavior of zeros of the corresponding OPs. As we learn from the work of Gonchar and Rakhmanov [12], greatly inspired by the earlier works of Stahl [31, 30], if there exists a contour $\Gamma_{*}\in\mathcal{T}$ for which the max-min problem (1.2) (when the supremum is taken over $K\in\mathcal{T}$) is attained, then the zeros of $p_{n}$ should accumulate, in the large degree limit, to the contour $\Gamma_{*}$. Furthermore, these zeros must also be distributed according to the equilibrium measure of $\Gamma_{*}$, that is, the probability measure for which the infimum in (1.2) over $\Gamma_{*}$ is attained.
@@
Our main result establishes the existence of the max-min contour for each of the classes $\mathcal{T}=\mathcal{T}_{\text{Herm}},\mathcal{T}=\mathcal{T}_{\text{Lag}},\mathcal{T}=\mathcal{T}_{\text{Jac}}$ as above, also for classes of contours obtained from unions of elements of those, and much more, allowing also for semiclassical external fields. Thus, with our work, we place many of the aforementioned works under the same framework, and showcase how one can leverage some techniques that have been developed and explored in the last years to more general versions of the Pólya-Chebotarev Problem.
@@ 2.1 Necessary concepts from potential theory
are called the logarithmic potential and the Cauchy transform of $\mu$, respectively. The logarithmic potential $\operatorname{U}^{\mu}$ is an harmonic function on $\mathbb{C}\setminus\operatorname{supp}\mu$, and under the additional condition
@@
For a closed set $F\subset\mathbb{C}$, we denote by $\mathcal{M}_{1}(F)$ the set of Borel probability measures supported on $F$ that satisfy (2.2) and have finite logarithmic energy. Given a function $\varphi$ defined on $F$, we denote by $\mathcal{M}_{1}^{\varphi}(F)$ the subset of measures $\mu\in\mathcal{M}_{1}(F)$ that also satisfy $\varphi\in L^{1}(\mu)$. Thus, for $\mu\in\mathcal{M}_{1}^{\varphi}(F)$ the (weighted logarithmic) energy
@@
is well defined, again with values in (,+]($\infty$,$\infty$]( - ∞ , + ∞ ].
@@
The measure $\mu_{F}$, when it exists, is called the equilibrium measure of $F$ in the external field $\varphi$ or simply the equilibrium measure of $F$ when $\varphi$ is understood. Let $\Gamma$ be a closed set in $\mathbb{C}$ that satisfies the growth condition
$\lim_{\begin{subarray}{c}z\to\infty\\ z\in\Gamma\end{subarray}}\varphi(z)-\log|z|=+\infty.$ (2.3)
Then the equilibrium measure $\mu(\varphi,\Gamma)$ uniquely exists [28] and is characterized by the Euler-Lagrange variational conditions: it is the unique measure in $\mathcal{M}_{1}^{\varphi}(\Gamma)$ for which there exists a constant $\ell$ satisfying
@@
In general, conditions (2.4) should be interpreted in the quasi-everywhere sense, but for us the external field will always be continuous except for finitely many points, so that (2.4) is true as stated.
When $\varphi\equiv 0$ and $F$ is a non-empty compact set, the class of measures $\mathcal{M}_{1}(F)=\mathcal{M}_{1}^{\varphi\equiv 0}(F)$ is non-empty and the problem just stated is well posed. In this case, if $\operatorname{I}(F)<+\infty$, then $\mu_{F}$ uniquely exists and is called the Robin measure of $F$, the value $\operatorname{I}(F)$ is called the Robin constant of $F$, and $\operatorname{cap}(F)=\mathrm{e}^{-\operatorname{I}(F)}$ is the (logarithmic) capacity of $F$.
Fix a finite set $\mathcal{C}\subset\mathbb{C}$. Still in this context of $\varphi\equiv 0$, the previously mentioned classical Pólya-Chebotarev Problem (1.1) was originally stated in the equivalent formulation of finding a compact and connected set $K_{*}$ with $\mathcal{C}\subset K_{*}$ and for which
@@
As mentioned, in the work we consider here we are interested in an extension of the classical Pólya-Chebotarev Problem (1.1), but now for the weighted energy $\operatorname{I}^{\varphi}$. We assume from now on that $\varphi$ is the real part of the sum of a rational function and a logarithm part. Without loss of generality, we place one of the poles of the rational function at $\infty$. Later on, when posing our extremal problem we will allow contours extending to $\infty$ along different sectors, and to avoid degenerate situations we assume that the pole at $\infty$ is of order at least $1$. In summary, throughout this paper we always assume that
@@
[…]
@@
The condition $N\geq 1$ ensures that $\infty$ is a pole of the rational part of $\Phi$, and also that condition (2.3) holds true along at least one nontrivial sector at $\infty$. The results we obtain could also be extended without the condition that $\rho\in\mathbb{C}\setminus\mathbb{R}$, as long as the points $w_{j}$ are not allowed to belong to the set of fixed points $\mathcal{C}$, but we opt for excluding this case from our analysis to avoid a separate technical analysis.
@@
and $\Phi^{\prime}$ is a rational function, which motivates us to call $\Phi$ a semiclassical external field. We also write
@@
that is, $\mathcal{Z}$ is the set of singularities of $\Phi$. The class of external fields we consider is now introduced, and we move on to the introduction of the class of contours of interest.
@@ 2.2 The class of admissible contours
In the classical problem (1.1), the class of sets consist of compact and connected sets containing $\mathcal{C}$. As shown by Stahl [31, 30], the set of minimal capacity $K_{*}$ is in fact a union of analytic arcs, and consequently the max-min problem (1.1) could well be posed on the more restrictive class of sets which are unions of contours containing $\mathcal{C}$, while still preserving its minimizer $K_{*}$.
In our extension, we will allow sets $\Gamma\in\mathcal{T}$ which are unions of contours with at least one of three distinguished features when compared with the classical Pólya-Chebotarev setup. The first feature is enforcing that prescribed subcollections of points in $\mathcal{C}$ must belong to the same connected components of a given $\Gamma\in\mathcal{T}$. The second feature is enforcing that prescribed subcollections of points of $\mathcal{C}$ to belong to the same connected components of $\Gamma\in\mathcal{T}$, and each of these connected components must extend to $\infty$ along prescribed angular sectors. The third feature is enforcing that each $\Gamma\in\mathcal{T}$ must have connected components stretching to $\infty$ along prescribed angular sectors. The canonical examples of connected components coming from each of these features are provided by contours in $\mathcal{T}_{\text{Jac}},\mathcal{T}_{\text{Lag}}$ and $\mathcal{T}_{\text{Herm}}$, respectively.
@@
The points in $\mathcal{C}$ will play the role of fixed points in the Pólya-Chebotarev problem. We also tacitly assume that
@@
$\varphi(z)=R^{N}|\alpha|\cos\left(N\theta+\arg(\alpha)\right)\left[1+\mathcal{O}(1/R)\right],\quad z\to+\infty$
@@
In particular, this implies that $\varphi(z)\to+\infty$ as $z\to\infty$ within $S_{j}\cap F$. Let $\Theta=\{\theta_{1}<\cdots<\theta_{N}\}$ be the set of the central angles of the admissible sectors as introduced in (2.11). We shall use $\theta_{j}$ to represent the sector $S_{j}$. We talk about partitions $\mathcal{P}(\Theta)$ of $\Theta$, which will be used when encoding the connection of different sectors at $\infty$ in our max-min problem. A partition $\mathcal{P}(\Theta)$ of $\Theta$ introduces an equivalence relation $\sim$ in the usual way: $\theta_{j}\sim\theta_{k}$ if both $\theta_{j}$ and $\theta_{k}$ belong to the same set in $\mathcal{P}(\Theta)$. We denote by $\mathcal{P}_{0}(\Theta)$ the subset of $\mathcal{P}(\Theta)$ obtained by removing all classes consisting of only one element. A partition $\mathcal{P}(\Theta)$ of $\Theta$ is said to be non-crossing if its associated equivalence relation $\sim$ satisfies the following property: if $\theta_{j}\sim\theta_{j^{\prime}}$ and $\theta_{k}\sim\theta_{k^{\prime}}$ with $j<k<j^{\prime}<k^{\prime}$, then necessarily $\theta_{j}\sim\theta_{k}$. Later on, we also talk about non-crossing partitions of subsets of $\Theta$.
The name non-crossing is better motivated when we represent this equivalence relation on the unit circle as follows. Represent each element $\theta$ of $\Theta$ on the unit circle as the point $(\cos(\theta),\sin(\theta))$ and connect consecutive elements of an equivalence class of the partition by a contour inside the circle, also connecting the first and last elements of the same class. A partition is then non-crossing if there exists a choice of these contours in such a way that different contours can only intersect if they connect elements of the same class. We refer the reader to Figures 0(a) and 0(b) for examples.
@@
In a similar manner, we will encode the connections of finite points through a partition $\mathcal{P}(\mathcal{C})$ of $\mathcal{C}$. As said earlier, we want to define classes of sets $\mathcal{T}$ for our Pólya-Chebotarev problem, which involves connecting finite points and $\infty$. In what follows, we talk about connecting points $c_{i},c_{j}\in\mathcal{C}$ through a set $\Gamma$, meaning that $c_{i}$ and $c_{j}$ belong to the same connected component of $\Gamma$. Likewise, we talk about connecting a point $c_{i}$ to an admissible sector $S_{k}$, meaning that $c_{i}$ belongs to a connected component of $\Gamma$ that stretches to $\infty$ along $S_{k}$. Finally, we also talk about connecting different sectors $S_{k}$ and $S_{l}$, meaning that there is a connected component of $\Gamma$ that stretches to $\infty$ in both sectors $S_{k}$ and $S_{l}$. The next step is to book-keep all such possible connections in a convenient notation, but before moving forward it is instructive to discuss the coming notions with an example.
@@
Fix a set of distinct points $\mathcal{C}=\{c_{1},\ldots,c_{6}\}$. Say that our external field determines exactly eight admissible sectors, the most trivial case being $\varphi(z)=\operatorname{Re}z^{8}$. We define now a class of sets $\mathcal{T}$ with the following features. First of all, a set $\Gamma\in\mathcal{T}$ must connect points in $\mathcal{C}$ with the following rules:
@@
  • Sectors $S_{4}$ and $S_{5}$ are connected.
Then the class $\{c_{1},c_{2},c_{3},c_{4}\}$ is connected to $S_{5}$, which is an information not contained in $\Psi$. Consider that
@@
While information above does not concerns any of the admissible sectors already present on function $\Psi$, it implies that $\{c_{1},c_{2},c_{3},c_{4}\}$ is connected to $S_{2}$ and $S_{5}$, which once again discredits $\Psi$. Finally, some redundancy can come just by the type of connections we would like to make, for example
@@
$\begin{split}\mathcal{P}(\mathcal{C})&=\{\{c_{1}\},\{c_{2},c_{3},c_{4}\},\{c_{5},c_{6}\}\}\\ \Psi(\{c_{1}\})&=\{\theta_{4}\},\quad\Psi(\{c_{2},c_{3},c_{4}\})=\emptyset,\quad\Psi(\{c_{5},c_{6}\})=\{\theta_{3}\}\\ \mathcal{P}(\Theta)&=\{\{\theta_{1}\},\{\theta_{2}\},\{\theta_{4}\},\{\theta_{3}\},\{\theta_{4}\},\{\theta_{5},\theta_{6}\},\{\theta_{7}\}\}\end{split}$ (2.14)
@@
Hence, in the previous example we determined $\mathcal{T}$ through the triple $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta_{\infty}))$. We now proceed to the discussion needed for the general definition. We book-keep subcollections of points of $\mathcal{C}$ that will be necessarily connected by contours in our Pólya-Chebotarev Problem by the classes of a given partition $\mathcal{P}(\mathcal{C})$ of $\mathcal{C}$, then two points are connected by a contour if, and only if, they are elements of the same partition.
@@
Both conditions are placed to avoid redundancies. Condition (2.15) means that if we have a fixed point not connected to any other fixed point, it must be connected to infinity through some admissible sector. We enforce this condition otherwise one could remove $w$ from the family of fixed points. Condition (2.16) means that we do not have different equivalence classes of finite points being enforced to connect to each other through infinity thorough the same sector, otherwise we should include then in the same class in $\mathcal{P}(\mathcal{C})$ from the very beginning.
@@
This solidify the codification nature of $\mathcal{P}(\mathcal{C})$ (see Example 2.1): by removing this condition one can create examples of $\Psi$ which enforces connections between fixed points in different classes of $\mathcal{P}(\mathcal{C})$.
@@
We say that a partition $\mathcal{P}(\Theta)$ is subordinated to the function $\Psi$ if $\mathcal{P}(\Theta)$ is a non-crossing partition of $\Theta$ with $\mathcal{P}_{\Psi}\subset\mathcal{P}(\Theta)$. The classes of $\mathcal{P}_{0}(\Theta)\setminus\mathcal{P}_{\Psi}$ codifies connections between admissible sectors not necessarily containing a fixed point, where two angles $\theta_{i}$ and $\theta_{j}$ are in the same class if, and only if, there is a contour stretching to infinity on $S_{j}$ and $S_{k}$. With the previous notations we have the necessary objects to define the class of admissible sets.

### v2

@@ 1 Introduction
A classical problem in Geometric Function Theory, known as the Pólya-Tchebotarev Problem, asks for finding a compact and connected set of minimal capacity that contains a given finite collection of points $\mathcal{C}\subset\mathbb{C}$ [28, 27]. In terms of the closely related notion of logarithmic energy, the Pólya-Tchebotarev Problem may be phrased as the question of finding a compact and connected set $K_{*}\subset\mathbb{C}$ containing $\mathcal{C}$, and for which the max-min (or sup-inf) problem
@@
The original form of the problem was presented by Pólya in 1929 [28], and solved by Lavrentieff [20, 19], Grötzch [15] and Goluzin [13] in the following years. Much later, motivated by questions on Padé approximants and rational approximation [38, 34, 26], Stahl established a more general version of Pólya-Tchebotarev problem [36, 35], in that it weakens the finiteness condition on $\mathcal{C}$, allowing it to be an arbitrary compact set with zero capacity.
Stahl’s ideas were rather influential in particular in the theory of orthogonal polynomials. In his developments, he also connected sets of minimal capacity with non-hermitian orthogonal polynomials, showing that in the large degree limit the zeros of such polynomials are described by sets of minimal capacity [37], see also [14]. Starting in the early 2000s, such ideas and results also found their way with applications to other areas of mathematics such as numerical analysis [12, 16, 9, 11, 25, 40, 2], random matrix theory [1, 3, 7, 4], integrable systems [8, 6, 17, 10], among others.
In many of these applications we just mentioned, the corresponding extremal problems that arise often are analogues of the original Pólya-Tchebotarev Problem, but now under the setting of weighted capacity, which is the main focus of our work here. We solve a weighted extension of the original Pólya-Tchebotarev Problem (1.1), obtained including an external field in the energy functional: for a given finite set $\mathcal{C}\subset\mathbb{C}$ and a suitable function $\varphi$ defined over $\mathbb{C}$, find a compact and connected set $K_{*}\subset\mathbb{C}$ containing $\mathcal{C}$, and for which the max-min problem
@@
When $\Phi$ is a constant/linear/quadratic real-valued polynomial and $\mathcal{T}=\mathcal{T}_{\text{Jac}}/\mathcal{T}_{\text{Lag}}/\mathcal{T}_{\text{Herm}}$, the polynomial $p_{n}$ reduces to (particular cases of rescaled) Jacobi/Laguerre/Hermite polynomial, respectively. For general complex-valued polynomial $\Phi$, the polynomial $p_{n}$, if it exists, is independent of the contour of orthogonality $\Gamma\in\mathcal{T}$ chosen, since the orthogonality condition (1.3) remains unchanged if we deform $\Gamma$ to a new contour $\widetilde{\Gamma}$ within the same class.
In each of the aforementioned cases, the flexibility in the choice of contours $\Gamma\in\mathcal{T}$ has impact in the behavior of zeros of the corresponding OPs. As we learn from the work of Gonchar and Rakhmanov [14], greatly inspired by the earlier works of Stahl [36, 35], if there exists a contour $\Gamma_{*}\in\mathcal{T}$ for which the max-min problem (1.2) (when the supremum is taken over $K\in\mathcal{T}$) is attained, then the zeros of $p_{n}$ should accumulate, in the large degree limit, to the contour $\Gamma_{*}$. Furthermore, these zeros must also be distributed according to the equilibrium measure of $\Gamma_{*}$, that is, the probability measure for which the infimum in (1.2) over $\Gamma_{*}$ is attained. Moreover, $\Gamma_{*}$ is uniquely determined up to a zero capacity set.
@@
Our main result establishes the existence of the max-min contour for each of the classes $\mathcal{T}=\mathcal{T}_{\text{Herm}},\mathcal{T}=\mathcal{T}_{\text{Lag}},\mathcal{T}=\mathcal{T}_{\text{Jac}}$ as above, also for classes of contours obtained from unions of elements of those, and much more, allowing also for semiclassical external fields. Thus, with our work, we place many of the aforementioned works under the same framework, and showcase how one can leverage some techniques that have been developed and explored in the last years to more general versions of the Pólya-Tchebotarev Problem.
@@ 2.1 Necessary concepts from potential theory
are called the logarithmic potential and the Cauchy transform of $\mu$, respectively. The logarithmic potential $\operatorname{U}^{\mu}$ is a harmonic function on $\mathbb{C}\setminus\operatorname{supp}\mu$, and under the additional condition
@@
For a set $F\subset\mathbb{C}$ which is closed, we denote by $\mathcal{M}_{1}(F)$ the set of Borel probability measures supported on $F$ that satisfy (2.2) and have finite logarithmic energy. Given a function $\varphi$ defined on $F$, we denote by $\mathcal{M}_{1}^{\varphi}(F)$ the subset of measures $\mu\in\mathcal{M}_{1}(F)$ that also satisfy $\varphi\in L^{1}(\mu)$. Thus, for $\mu\in\mathcal{M}_{1}^{\varphi}(F)$ the (weighted logarithmic) energy
@@
is well defined, and takes values in $\mathbb{R}$.
@@
The measure $\mu_{F}$, when it exists, is called the equilibrium measure of $F$ in the external field $\varphi$ or simply the equilibrium measure of $F$ when $\varphi$ is understood.
Let $\Gamma$ be a closed set in $\mathbb{C}$ that satisfies the growth condition
$\lim_{\begin{subarray}{c}z\to\infty\\ z\in\Gamma\end{subarray}}\left[\varphi(z)-\log|z|\right]=+\infty,$ (2.3)
and let $\mathcal{Z}$ be the set of discontinuities of $\varphi$. For us the $\varphi$ shall always be such that $\mathcal{Z}$ is finite. Then the equilibrium measure $\mu(\varphi,\Gamma)$ uniquely exists [33], provided that $\Gamma$ has positive capacity, and is characterized by the Euler-Lagrange variational conditions: it is the unique measure in $\mathcal{M}_{1}^{\varphi}(\Gamma)$ for which there exists a constant $\ell$ satisfying
@@
In general, conditions (2.4) should be interpreted in the quasi-everywhere sense, but whenever $\mathcal{Z}$ is finite and $\Gamma$ is a finite union of continua (compact sets with more than one point), they hold as stated.
When $\varphi\equiv 0$ and $F$ is a compact set with positive capacity, the class of measures $\mathcal{M}_{1}(F)=\mathcal{M}_{1}^{\varphi\equiv 0}(F)$ is non-empty and the problem just stated is well posed. In this case $\mu_{F}$ uniquely exists and is called the Robin measure of $F$, the value $\operatorname{I}(F)$ is called the Robin constant of $F$, and $\operatorname{cap}(F)=\mathrm{e}^{-\operatorname{I}(F)}$ is the (logarithmic) capacity of $F$.
Fix a finite set $\mathcal{C}\subset\mathbb{C}$. Still in this context of $\varphi\equiv 0$, the previously mentioned classical Pólya-Tchebotarev Problem (1.1) was originally stated in the equivalent formulation of finding a compact and connected set $K_{*}$ with $\mathcal{C}\subset K_{*}$ and for which
@@
As mentioned, in the work we consider here we are interested in an extension of the classical Pólya-Tchebotarev Problem (1.1), but now for the weighted energy $\operatorname{I}^{\varphi}$. We assume from now on that $\varphi$ is the real part of the sum of a rational function and a logarithm part. Without loss of generality, we place one of the poles of the rational function at $\infty$. Later on, when posing our extremal problem we will allow contours extending to $\infty$ along different sectors, and to avoid degenerate situations we assume that the pole at $\infty$ is of order at least $1$. In summary, throughout this paper we always assume that
@@
$Q(z)\mathrel{\vbox{\hbox{\scriptsize.}\hbox{\scriptsize.}}}=\prod_{j=1}^{n_{Q}}(z-z_{j})^{m_{j}},\quad m_{j}>0,\quad j=1,\ldots,n_{Q},$ (2.6)
@@
Lastly, we assume that all $z_{j}$ and $w_{j}$ are pairwise distinct.
The condition $N\geq 1$ ensures that $\infty$ is a pole of the rational part of $\Phi$, and also that condition (2.3) holds true along at least one nontrivial sector at $\infty$.
@@
and $\Phi^{\prime}$ is a rational function, which motivates us to call $\Phi$ a semiclassical external field.
We also write
@@
that is, $\mathcal{Z}$ is the set of singularities of $\Phi$.
The class of external fields we consider is now introduced, and we move on to the introduction of the class of contours of interest.
@@ 2.2 The class of admissible contours
In the classical problem (1.1), the class of sets consist of compact and connected sets containing $\mathcal{C}$. As shown by Stahl [36, 35], the set of minimal capacity $K_{*}$ is in fact a union of analytic arcs, and consequently the max-min problem (1.1) could as well be posed on the more restrictive class of sets which are unions of contours containing $\mathcal{C}$, while still preserving its minimizer $K_{*}$.
In our extension, we will allow sets $\Gamma\in\mathcal{T}$ which are unions of contours with at least one of three distinguished features when compared with the classical Pólya-Tchebotarev setup. The first feature is enforcing that prescribed subcollections of points in $\mathcal{C}$ must belong to the same connected components of a given $\Gamma\in\mathcal{T}$. The second feature is enforcing that prescribed subcollections of points of $\mathcal{C}$ to belong to the same connected components of $\Gamma\in\mathcal{T}$, and each of these connected components must extend to $\infty$ along prescribed angular sectors. The third feature is enforcing that each $\Gamma\in\mathcal{T}$ must have connected components stretching to $\infty$ along prescribed angular sectors. The canonical examples of connected components coming from each of these features are provided by contours in $\mathcal{T}_{\text{Jac}},\mathcal{T}_{\text{Lag}}$ and $\mathcal{T}_{\text{Herm}}$, respectively.
@@
[…]
We also throughout the paper assume that
@@
$\varphi(z)=R^{N}|\alpha|\cos\left(N\theta+\arg(\alpha)\right)\left[1+o(1)\right],\quad z\to+\infty$
@@
In particular, this implies that $\varphi(z)\to+\infty$ as $z\to\infty$ within $S_{j}\cap F$.
Let $\Theta=\{\theta_{1}<\cdots<\theta_{N}\}$ be the set of the central angles of the admissible sectors as introduced in (2.11). We shall use $\theta_{j}$ to represent the sector $S_{j}$. We talk about partitions $\mathcal{P}(\Theta)$ of $\Theta$, which will be used when encoding the connection of different sectors at $\infty$ in our max-min problem. A partition $\mathcal{P}(\Theta)$ of $\Theta$ introduces an equivalence relation $\sim$ in the usual way: $\theta_{j}\sim\theta_{k}$ if both $\theta_{j}$ and $\theta_{k}$ belong to the same set in $\mathcal{P}(\Theta)$. We denote by $\mathcal{P}_{0}(\Theta)$ the subset of $\mathcal{P}(\Theta)$ obtained by removing all classes consisting of only one element.
A partition $\mathcal{P}(\Theta)$ of $\Theta$ is said to be non-crossing if its associated equivalence relation $\sim$ satisfies the following property: if $\theta_{j}\sim\theta_{j^{\prime}}$ and $\theta_{k}\sim\theta_{k^{\prime}}$ with $j<k<j^{\prime}<k^{\prime}$, then necessarily $\theta_{j}\sim\theta_{k}$. Later on, we also talk about non-crossing partitions of subsets of $\Theta$.
The name non-crossing is better motivated when we represent this equivalence relation on the unit circle as follows. Represent each element $\theta$ of $\Theta$ on the unit circle as the point $(\cos(\theta),\sin(\theta))$ and connect consecutive elements of an equivalence class of the partition by a contour inside the circle, also connecting the first and last elements of the same class. A partition is then non-crossing if there exists a choice of these contours in such a way that different contours can only intersect if they connect elements of the same class. We refer the reader to Figures 1(a) and 1(b) for examples.
@@
In a similar manner, we will encode the connections of finite points through a partition $\mathcal{P}(\mathcal{C})$ of $\mathcal{C}$. As said earlier, we want to define classes of sets $\mathcal{T}$ for our Pólya-Tchebotarev problem, which involves connecting finite points and $\infty$. In what follows, we talk about connecting points $c_{i},c_{j}\in\mathcal{C}$ through a set $\Gamma$, meaning that $c_{i}$ and $c_{j}$ belong to the same connected component of $\Gamma$. Likewise, we talk about connecting a point $c_{i}$ to an admissible sector $S_{k}$, meaning that $c_{i}$ belongs to a connected component of $\Gamma$ that stretches to $\infty$ along $S_{k}$. Finally, we also talk about connecting different sectors $S_{k}$ and $S_{l}$, meaning that there is a connected component of $\Gamma$ that stretches to $\infty$ in both sectors $S_{k}$ and $S_{l}$.
The next step is to book-keep all such possible connections in a convenient notation, but before moving forward it is instructive to discuss the coming notions with an example.
@@
Fix a set of distinct points $\mathcal{C}=\{c_{1},\ldots,c_{6}\}$. Say that our external field determines exactly eight admissible sectors, the most trivial case being $\varphi(z)=\operatorname{Re}z^{8}$. We define now a class of sets $\mathcal{T}$ with the following features.
First of all, a set $\Gamma\in\mathcal{T}$ must connect points in $\mathcal{C}$ with the following rules:
@@
  • Sectors $S_{1}$ and $S_{5}$ are connected.
Then the class $\{c_{1},c_{2},c_{3},c_{4}\}$ is connected to $S_{5}$ (through infinity), which is an information not contained in $\Psi$. Consider that
@@
While information above does not concern any of the admissible sectors already present on function $\Psi$, it implies that $\{c_{1},c_{2},c_{3},c_{4}\}$ is connected to $S_{2}$ and $S_{5}$, which once again discredits $\Psi$. Finally, some redundancy can come just by the type of connections we would like to make, for example
@@
Assuming (2.12) and (2.13), we have the correct levels of codification on the sets $\mathcal{P}(\mathcal{C})$, $\mathcal{P}_{\Psi}$ and $\mathcal{P}(\Theta)$. For example, if $\mathcal{C}$ has $6$ elements and $\Theta$ has seven, putting
$\begin{split}\mathcal{P}(\mathcal{C})&=\{\{c_{1}\},\{c_{2},c_{3},c_{4}\},\{c_{5},c_{6}\}\}\\ \Psi(\{c_{1}\})&=\{\theta_{4}\},\quad\Psi(\{c_{2},c_{3},c_{4}\})=\emptyset,\quad\Psi(\{c_{5},c_{6}\})=\{\theta_{3}\}\\ \mathcal{P}(\Theta)&=\{\{\theta_{1}\},\{\theta_{2}\},\{\theta_{3}\},\{\theta_{4}\},\{\theta_{5},\theta_{6}\},\{\theta_{7}\}\}\end{split}$ (2.14)
@@
Hence, in the previous example we determined $\mathcal{T}$ through the triple $(\mathcal{P}(\mathcal{C}),\Psi,\mathcal{P}(\Theta_{\infty}))$. We now proceed to the discussion needed for the general definition.
We book-keep subcollections of points of $\mathcal{C}$ that will be necessarily connected by contours in our Pólya-Tchebotarev Problem by the classes of a given partition $\mathcal{P}(\mathcal{C})$ of $\mathcal{C}$, then two points are connected by a contour if, and only if, they are elements of the same partition.
@@
Both conditions are placed to avoid redundancies. Condition (2.15) means that if we have a fixed point not connected to any other fixed point, it must be connected to infinity through some admissible sector. We enforce this condition otherwise one could remove $w$ from the family of fixed points. Condition (2.16) means that we do not have different equivalence classes of finite points being enforced to connect to each other through infinity through the same sector, otherwise we should include then in the same class in $\mathcal{P}(\mathcal{C})$ from the very beginning.
@@
This solidifies the codification nature of $\mathcal{P}(\mathcal{C})$ (see Example 2.1): by removing this condition one can create examples of $\Psi$ which enforces connections between fixed points in different classes of $\mathcal{P}(\mathcal{C})$.
@@
We say that a partition $\mathcal{P}(\Theta)$ is subordinated to the function $\Psi$ if $\mathcal{P}(\Theta)$ is a non-crossing partition of $\Theta$ with $\mathcal{P}_{\Psi}\subset\mathcal{P}(\Theta)$. The classes of $\mathcal{P}_{0}(\Theta)\setminus\mathcal{P}_{\Psi}$ codify connections between admissible sectors not necessarily containing a fixed point, where two angles $\theta_{i}$ and $\theta_{j}$ are in the same class if, and only if, there is a contour stretching to infinity on $S_{j}$ and $S_{k}$.
With the previous notations we have the necessary objects to define the class of admissible sets.